Two stock cars race on a 5km circuit and pass a marked simultaneously. How long will it take the faster car to gain one lap on the slower if their average speeds are:
120km/h and 100km/h?
It says on the answer sheet the answer is 15 minutes, I'd much appreciate it if someone would teach me.

Answer by Alan3354(69443)

Two stock cars race on a 5km circuit and pass a marked simultaneously. How long will it take the faster car to gain one lap on the slower if their average speeds are:
120km/h and 100km/h?
-------------
One lap = 5 km
The difference in speed is 20 km/hr
t = d/r = 5/20 = 1/4 hr = 15 minutes